# The Effect of Omitted Variables on

### v3

@@ 3.2 Incorporating Magnitude Restrictions on the Omitted Variable Bias
A5 says that the magnitude of the omitted variable bias is no larger than $M$. This assumption can be viewed as formalizing Oster’s (2019) statement that researchers may be “willing to assume that the bias is fairly small” (page 194). The value $M$ should be chosen to reflect the researcher’s beliefs about the largest possible bias in their baseline estimand $\beta_{\text{med}}$. For example, consider a regression of log earnings $Y$ on years of schooling $X$ and other observed covariates $W_{1}$. Suppose our estimate of $\beta_{\text{med}}$ is 0.1. Then $M=1$ would say that the true coefficient on schooling—as obtained after adjusting for the unobserved confounders—is at most 1.1. This value is far outside the range of most estimates (e.g., Card 2001 table II) and so in this application researchers would likely view A5 with $M=1$ as a plausible assumption.
@@
$\beta_{\text{long}}\in\big{[}\beta_{\text{med}}-p|\beta_{\text{med}}|,\beta_{\text{med}}+p|\beta_{\text{med}}|\big{]}.$
Every element of this set has the same sign if and only if $p<1$. For this reason, since we are interested in robustness to sign changes, $p$ should be chosen to be larger than 1. In the returns to schooling example with $\beta_{\text{med}}=0.1$, $M=1$ is equivalent to the choice $p=10$. We recommend that researchers consider a range of magnitude restrictions, as we illustrate empirically in section 6. For example, in section 6 we consider $p=2$, $3$, and $10$.
@@
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1A5 hold. Suppose $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$ is known. Recall that A2 specifies that $\delta\in\mathbb{R}$ is known. Then the identified set for $\beta_{\text{long}}$ is
$\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M)\coloneqq\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})\cap[\beta_{\text{med}}-M,\beta_{\text{med}}+M].$
@@
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1, A3, A4, and A6 hold. Suppose $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$ is known. Then the identified set for $\beta_{\text{long}}$ is
¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})\coloneqq\bigcup_{|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2}).$
Suppose A5 also holds. Then the identified set for $\beta_{\text{long}}$ is
¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2},M)\coloneqq\bigcup_{|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M).$
@@ 5.1 $\delta$ is a Double Ratio
This property that $\delta$ is a ratio of ratios leads to some counter-intuitive interpretations of the magnitude of $\delta$. For example, suppose $\gamma_{2,\text{long}}=c\cdot\gamma_{1,\text{long}}$ and $\pi_{2}=c\cdot\pi_{1}$ for $c=10$. Since both $W_{1}$ and $W_{2}$ are normalized to have unit variance, this value of $c$ indicates that $W_{2}$ has an association with $X$ (and $Y$) that is 10 times stronger than that of $W_{1}$ with $X$ (and $Y$). However, the implied value of $\delta$ is 1, which is traditionally interpreted as meaning that there is “equal selection” based on $W_{2}$ and on $W_{1}$. This value of $\delta$ is unchanged if $c=1/10$, in which case $W_{2}$’s association with $X$ and $Y$ is 10 times weaker than $W_{1}$’s association with those same variables. In fact, the value of $\delta$ is equal to 1 for all $c\neq 0$ in this example. As Cinelli and Hazlett (2020) point out, the interpretation of $\delta$ in some papers is at odds with its double ratio property. For example Oster (2019, page 192) states that $\delta=1$ implies that “the unobservable[s] and observables are equally related to the treatment.” Their paper recommends continuing to use $R_{\text{long}}^{2}$ as one of the sensitivity parameters, but that $\delta$ should be replaced with RXW2
W1
2
@@
[…]
, where RXW2
W1
2
[…]
is the unknown $R$-squared in a regression of $X$ on $W_{2}$, after partialling out $W_{1}$.
5.2 Exogeneity of the Controls
All of the sensitivity analyses we discussed in sections 24 maintain assumption A4. That assumption requires all observed variables to be uncorrelated with all omitted variables. In many applications, this assumption would be viewed as strong—most observed variables are themselves correlated and it is unlikely that the omitted variables happen to be uncorrelated with all of the observed covariates, especially when many observed covariates are included in the analysis. Oster argues this assumption can be relaxed by redefining $W_{2}$ to be $W_{2}^{\perp W_{1}}=W_{2}-\operatorname*{cov}(W_{2},W_{1})\operatorname*{var}(W_{1})^{-1}W_{1}$, which is uncorrelated with $W_{1}$ by definition. However, as Diegert et al. (2022) argue, this redefinition of $W_{2}$ substantively changes the interpretation of $\delta$. The interpretation changes because the replacement of $W_{2}$ by $W_{2}^{\perp W_{1}}$ requires replacing the coefficients $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$ with $(\widetilde{\gamma}_{1},\gamma_{2,\text{long}})$ where $\widetilde{\gamma}_{1}=\gamma_{1,\text{long}}+\gamma_{2,\text{long}}^{\prime}\operatorname*{cov}(W_{2},W_{1})\operatorname*{var}(W_{1})^{-1}$. The redefined $\delta$ is thus not only a function of the associations between $(W_{1},W_{2})$ and $X$ and $Y$, as discussed above, but also of the unknown covariance between $W_{2}$ and $W_{1}$. Concretely, the redefined $\delta$ is
$\delta=\frac{\pi_{2}}{\pi_{1}}\Bigg{/}\frac{\gamma_{2,\text{long}}}{\gamma_{1,\text{long}}+\gamma_{2,\text{long}}\frac{\operatorname*{cov}(W_{2},W_{1})}{\operatorname*{var}(W_{1})}}.$
This concern can be addressed by using an alternative parameterization of the omitted variable bias formula, such as those in Cinelli and Hazlett (2020) or Diegert et al. (2022). The parameterization in Cinelli and Hazlett (2020) continues to use $R_{\text{long}}^{2}$ but replaces $\delta$ with RXW2
W1
2
[…]
2
[…]
(see section 5.1 above). The parameterization in Diegert et al. (2022) replaces both $\delta$ and $R_{\text{long}}^{2}$ with $r_{X}^{2}\coloneqq\operatorname*{var}(\pi_{2}^{\prime}W_{2})/\operatorname*{var}(\pi_{1}^{\prime}W_{1})$ and $r_{Y}^{2}\coloneqq\operatorname*{var}(\gamma_{2,\text{long}}^{\prime}W_{2})/\operatorname*{var}(\gamma_{1,\text{long}}^{\prime}W_{1})$, which are measures of the relative importance of $W_{2}$ and $W_{1}$ in the selection and outcome equation, respectively.
@@ 6 Empirical Application: Social Capital and the Rise of the Nazi Party
Their paper has a variety of analyses. We focus on their main results, which use a selection on observables identification strategy paired with linear models. That is, they estimate linear regressions of $Y$ on $X$, an intercept, and additional observed covariates. To assess the robustness of those results to omitted variables, they applied Oster (2019) in their subsection 4.D, with additional details in their appendix G. We re-examine that analysis in light of our theoretical results in section 3. See appendix LABEL:sec:extraEmpirical for extra details and results for this empirical application.
@@
Table 6 shows the main breakdown results. This table replicates and extends the bottom half of Panel B in Table A.27 of appendix G in Satyanath et al. (2017). The three rows of Table 6 correspond to the three different versions of treatment. All of the cell entries are different kinds of breakdown points. First consider columns (1) and (2). These show estimates of Oster’s explain away breakdown point. We present estimates for two different choices of $R_{\text{long}}^{2}$, either 1 or Oster’s rule of thumb $1.3\widehat{R}_{\text{med}}^{2}$. Column (1) shows the values reported by Satyanath et al. (2017). As described in appendix LABEL:sec:commonMistake, these values are incorrect. Column (2) reports the correct values. If we judged robustness by comparing the magnitude of these values of the explain away breakdown point to 1, then all of the results would be deemed robust except one, civic associations with $R_{\text{long}}^{2}=1$, which gives an explain away breakdown point of 0.98.
@@
Corollary 2, ^$\widehat{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1$ $(-\infty,-0.0855]\cup[0.0432,1.8947]$
Corollary 2, ^$\widehat{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1.05$ $(-\infty,-0.0798]\cup[0.0415,\infty)$
@@
Identified Set Plots
@@ 7.1 Data
The second meta-analysis is based on papers used in the meta-analysis of Oster (2019, sections 2 and 5). This dataset has 141 regressions from 55 different papers that were published in either a top 5 economics journal or the AEJ: Applied Economics between 2008 and 2013, and which reported results from coefficient stability analyses. See Oster (2019) for more details on the construction of this dataset. Since this dataset already only contains only a handful of regressions per paper, we did not further restrict the sample. This second dataset is useful both because it provides a larger sample of regressions and also because it the inclusion criteria are qualitatively different, since none of these papers implemented Oster’s methods because they preceded her paper.
@@ 7.2 Results
First consider the sample in Panel A. As shown in Theorem 2, the sign change breakdown point without any magnitude restrictions is never larger than 1. This result can be seen in the last line of all panels. For the rule of thumb choice of $R_{\text{long}}^{2}$, most of these breakdown points exactly equal 1. For $R_{\text{long}}^{2}=1$, more than half of the regressions are no longer robust when using $\delta=1$ as the robustness threshold point. In practice, researchers almost always go beyond binary robust / not-robust conclusions and also report the magnitude of the explain away breakdown point, interpreting larger magnitudes as evidence of greater robustness. The second row of each sub-panel shows that even with a reasonably strong restriction on the magnitude of OVB, the sign change breakdown points are substantially smaller than the explain away breakdown points. For example, the median is cut in half. Near the top of the distribution, the 90th percentile shrinks 87% from 13.22 down to 1.66 for the rule of thumb $R_{\text{long}}^{2}$, and similarly for $R_{\text{long}}^{2}=1$. The results for Panel B are qualitatively similar: The sign change breakdown points are substantially smaller in magnitude than the explain away breakdown points. Overall, these results show that in most empirical applications, common conclusions like
“selection on unobservables would have to be at least 4.1 times stronger than selection on observables to explain away the relationship” (Squicciarini 2020, AER, page 20)
@@
$R^{2}_{\text{long}}$ $\delta_{\text{hypo}}$   10% 25% 50% 75% 90%
@@
$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ $\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ 1.0 1.0 1.0 1.3 67.2
2 1.0 1.4 3.4 6.9 67.2
1.05 2.0 9.2 25.2 54.0 198.9
1 2.1 709.8 $\infty$ $\infty$ $\infty$
@@
$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ $\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ 1.0 1.0 1.0 2.7 11.4
2 1.0 3.3 12.8 58.6 116.0
1.05 3.7 16.8 70.7 255.0 771.6
1 5.3 59.6 1978.6 $\infty$ $\infty$
[…]
2 1.0 2.1 10.8 45.8 211.8
1.05 1.5 16.3 50.5 192.3 760.0
1 1.6 32.1 216.4 $\infty$ $\infty$
[…]
  • Note: Both samples only include results considered robust according to standard practice (the explain away breakdown point is weakly larger than 1, where we use the value of $R_{\text{long}}^{2}$ specified in the corresponding panel). Panel A: Top: $N=29$ regressions from 11 papers. Bottom: $N=11$ regressions from 6 papers. Panel B: $N=111$ regressions from 47 papers. Bottom: $N=58$ regressions from 31 papers. The value $\infty$ means our optimizer to compute $\widehat{p}_{i}^{\text{bp}}$ stopped at the preset upper bound of 1 million, which can be interpreted as effectively infinite. For $\delta_{\text{hypo}}=2$ and 1.05, we actually use $\min\{2,\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})\}$ and $\min\{1.05,\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})\}$ to ensure that the distributions in each sub-panel are stochastically ranked from top to bottom.
@@
To make magnitudes of OVB comparable across different regressions, we work with magnitudes relative to estimates of the baseline estimand $\beta_{\text{med}}$. Specifically, for the $i$th regression, let $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ be the largest magnitude of OVB (defined as a proportion of the baseline estimate) such that the sign change breakdown point is at least as large as $\delta_{\text{hypo}}$. Setting $\delta_{\text{hypo}}=\delta_{i}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2})$ gives the answer to our question above—when can the explain away breakdown point be interpreted as a sign change breakdown point? [^4] Formally, we define $p_{i}^{\text{bp}}(\delta_{\text{hypo}})\coloneqq\sup\{p\geq 1:|\delta_{i}^{\text{bp},\text{sign}}(R_{\text{long}}^{2},p\cdot|\beta_{\text{med},i}|)|\geq|\delta_{\text{hypo}}|\}$. Denote its estimator by $\widehat{p}_{i}^{\text{bp}}(\delta_{\text{hypo}})\coloneqq\sup\{p\geq 1:|\widehat{\delta}_{i}^{\text{bp},\text{sign}}(R_{\text{long}}^{2},p\cdot|\widehat{\beta}_{\text{med},i}|)|\geq|\delta_{\text{hypo}}|\}$.
@@
The first row of all four sub-panels in table 7.2 give the distribution of $\widehat{p}_{i}^{\text{bp}}(\widehat{\delta}^{\text{bp},\text{explain away}}(R_{\text{long}}^{2}))$ across regressions $i$ in our meta-analysis samples, and for different choices of $R_{\text{long}}^{2}$. Here we see that, across both samples and both choices of $R_{\text{long}}^{2}$, more than 50% of regressions require that the sign of the parameter of interest $\beta_{\text{long}}$ must be known a priori if we want to interpret the explain away breakdown point as a sign change breakdown point. That follows since—recall from section 3.2—the choice $M=\beta_{\text{med}}$ (equivalently, $p=1$) implies that the magnitude restriction A5 is equivalent to $\beta_{\text{long}}\in[0,2\beta_{\text{med}}]$ (when the baseline estimand is positive; the negative case is analogous). If the sign of $\beta_{\text{long}}$ was known a priori, however, then there would be no need to perform a sensitivity analysis to assess whether omitted variables could flip the sign of one’s baseline results; they cannot by assumption.
@@
For regressions above the 50th percentile, the implicit magnitude restrictions is typically larger than $p=1$, but is often still quite restrictive. For example, the 75th percentile in panel A is 1.3 under the rule of thumb $R_{\text{long}}^{2}$ choice. To interpret this, suppose $\widehat{\beta}_{\text{med}}=2$. Then $p=1.3$ corresponds to the assumption that $\beta_{\text{long}}\in[-0.6,4.6]$.
So far we have focused on the magnitude restriction required to interpret the explain away breakdown point as a sign change breakdown point. We can use $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ to answer several related questions as well. For example, what is the weakest possible restriction on the magnitude of OVB to ensure that the sign change breakdown point is at least 1? $p_{i}(\delta_{\text{hypo}})$ with $\delta_{\text{hypo}}=1$ answers this question. The bottom row of all sub-panels shows the distribution of our estimates of these magnitudes. Here we see that even for this very weak conclusion, at least 10% of regressions still require substantial magnitude restrictions. Most papers, however, require a fairly weak, or essentially no magnitude restriction in order to obtain a sign change breakdown point equal to or larger than 1. However, because of the asymptote at $\delta=1$, these $p_{i}^{\text{bp}}$ values may not be informative about values for $\delta$ cutoffs slightly above 1. So the middle rows of each sub-panel show estimates of $p_{i}^{\text{bp}}(\delta_{\text{hypo}})$ for values of $\delta_{\text{hypo}}$ close to, but slightly above 1. In these middle rows we see that obtaining sign change breakdown points which are meaningfully larger than 1 often requires substantial magnitude restrictions.

### v4

@@ 3.2 Incorporating Magnitude Restrictions on the Omitted Variable Bias
A5 says that the magnitude of the omitted variable bias is no larger than $M$. This assumption can be viewed as formalizing Oster’s (2019) statement that researchers may be “willing to assume that the bias is fairly small” (page 194). The value $M$ should be chosen to reflect the researcher’s beliefs about the largest possible bias in their baseline estimand $\beta_{\text{med}}$. For example, consider a regression of log earnings $Y$ on years of schooling $X$ and other observed covariates $W_{1}$. Suppose our estimate of $\beta_{\text{med}}$ is 0.1. Then $M=1$ would say that the true coefficient on schooling—as obtained after adjusting for the unobserved confounders—lies between $-0.9$ and 1.1. This interval includes most estimates (e.g., Card 2001 Table II) so in this application researchers would likely view A5 with $M=1$ as a plausible assumption.
@@
$\beta_{\text{long}}\in\big{[}\beta_{\text{med}}-p\cdot|\beta_{\text{med}}|,\beta_{\text{med}}+p\cdot|\beta_{\text{med}}|\big{]}.$
All elements of this set have the same sign if $p<1$. For this reason, since we are interested in robustness to sign changes, $p$ should be chosen to be larger than 1. In the returns to schooling example with $\beta_{\text{med}}=0.1$, $M=1$ is equivalent to the choice $p=10$. We recommend that researchers consider a range of magnitude restrictions, as we illustrate empirically in Section 6. For example, in Section 6 we consider $p=2$, $3$, and $10$.
@@
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1A5 hold. Suppose $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$ is known. Recall that A2 specifies that $\delta\in\mathbb{R}$ is known. Denote the identified set for $\beta_{\text{long}}$ under these assumptions by $\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M)$. Then
$\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M)=\mathcal{B}_{I}(\delta,R_{\text{long}}^{2})\cap[\beta_{\text{med}}-M,\beta_{\text{med}}+M].$
@@
Suppose the joint distribution of $(Y,X,W_{1})$ is known. Suppose A1, A3, A4, and A6 hold. Suppose $R_{\text{long}}^{2}\in(R_{\text{med}}^{2},1]$ is known. Denote the identified set for $\beta_{\text{long}}$ under these assumptions by ¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})$. Then
¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2})=\bigcup_{\delta:|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2}).$
Suppose A5 also holds. Denote the identified set for $\beta_{\text{long}}$ under this as well as the earlier assumptions by ¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2},M)$. Then
¯$\overline{\mathcal{B}}_{I}(\bar{\delta},R_{\text{long}}^{2},M)=\bigcup_{\delta:|\delta|\leq\bar{\delta}}\mathcal{B}_{I}(\delta,R_{\text{long}}^{2},M).$
@@ 5.1 $\delta$ is a Double Ratio
5.2 Exogeneity of the Controls
All of the sensitivity analyses we discussed in sections 24 maintain Assumption A4. That assumption requires all observed variables to be uncorrelated with all omitted variables. In many applications, this assumption would be viewed as strong—most observed variables are themselves correlated and it is unlikely that the omitted variables happen to be uncorrelated with all of the observed covariates, especially when many observed covariates are included in the analysis. Oster argues this assumption can be relaxed by redefining $W_{2}$ to be $W_{2}^{\perp W_{1}}:=W_{2}-\operatorname*{cov}(W_{2},W_{1})\operatorname*{var}(W_{1})^{-1}W_{1}$, which is uncorrelated with $W_{1}$ by definition. However, as Diegert et al. (2025b) argue, this redefinition of $W_{2}$ substantively changes the interpretation of $\delta$. The interpretation changes because the replacement of $W_{2}$ by $W_{2}^{\perp W_{1}}$ requires replacing the coefficients $(\gamma_{1,\text{long}},\gamma_{2,\text{long}})$ with $(\widetilde{\gamma}_{1},\gamma_{2,\text{long}})$ where $\widetilde{\gamma}_{1}:=\gamma_{1,\text{long}}+\operatorname*{var}(W_{1})^{-1}\operatorname*{cov}(W_{1},W_{2})\gamma_{2,\text{long}}$. The redefined $\delta$ is thus not only a function of the associations between $(W_{1},W_{2})$ and $X$ and $Y$, as discussed above, but also of the unknown covariance between $W_{2}$ and $W_{1}$. Concretely, for scalar $W_{1}$ and $W_{2}$, the redefined $\delta$ is
$\widetilde{\delta}$ $\coloneqq\frac{\operatorname*{cov}(X,\gamma_{2,\text{long}}W_{2}^{\perp W_{1}})}{\operatorname*{var}(\gamma_{2,\text{long}}W_{2}^{\perp W_{1}})}\Bigg{/}\frac{\operatorname*{cov}(X,\widetilde{\gamma}_{1}W_{1})}{\operatorname*{var}(\widetilde{\gamma}_{1}W_{1})}$
$=\frac{\pi_{2}}{\pi_{1}+\pi_{2}\frac{\operatorname*{cov}(W_{2},W_{1})}{\operatorname*{var}(W_{1})}}\Bigg{/}\frac{\gamma_{2,\text{long}}}{\gamma_{1,\text{long}}+\gamma_{2,\text{long}}\frac{\operatorname*{cov}(W_{2},W_{1})}{\operatorname*{var}(W_{1})}}.$
@@ 6 Empirical Application: Social Capital and the Rise of the Nazi Party
Their paper has a variety of analyses. We focus on their main results, which use a selection on observables identification strategy paired with linear models. That is, they estimate linear regressions of $Y$ on $X$, an intercept, and additional observed covariates. To assess the robustness of those results to omitted variables, they applied Oster (2019) in their Subsection 4.D, with additional details in their Appendix G. We re-examine that analysis in light of our theoretical results in Section 3. See Appendix LABEL:sec:extraEmpirical_details for extra details and results for this empirical application.
@@
Table 6 shows the main breakdown results. This table replicates and extends the bottom half of Panel B in Table A.27 of Appendix G in Satyanath et al. (2017). The three panels of Table 6 correspond to the three different versions of treatment. All of the cell entries are different kinds of breakdown points. First consider columns (1) and (2). These show estimates of Oster’s explain away breakdown point. We present estimates for two different choices of $R_{\text{long}}^{2}$, either 1 or Oster’s rule of thumb $1.3\widehat{R}_{\text{med}}^{2}$. Column (1) shows the values reported by Satyanath et al. (2017). As described in Appendix LABEL:sec:commonMistake, these values are incorrect. Column (2) reports the correct values. If we judged robustness by comparing the magnitude of these values of the explain away breakdown point to 1, then all of the results would be deemed robust except one, civic associations with $R_{\text{long}}^{2}=1$, which gives an explain away breakdown point of 0.98.
@@
Corollary 2, ¯^$\widehat{\overline{\mathcal{B}}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1$ $(-\infty,-0.0855]\cup[0.0432,1.8947]$
Corollary 2, ¯^$\widehat{\overline{\mathcal{B}}}_{I}(\bar{\delta},R_{\text{long}}^{2})$ $|\delta|\leq\bar{\delta}=1.05$ $(-\infty,-0.0798]\cup[0.0415,\infty)$
@@
Identified Set Plots and Main Empirical Recommendations
@@
Researchers may also want to summarize the information shown in plots like Figure 2 of the cumulative identified set. We recommend two ways in which researchers can do this, both of which modify and extend current practice as documented in Appendix LABEL:sec:inPractice.
  1. 1.
    First, in lieu of computing bias adjusted estimates based on equation (7), researchers can estimate the cumulative identified set ¯^$\widehat{\overline{\mathcal{B}}}_{I}(\overline{\delta},R_{\text{long}}^{2})$ for a reference value of $\overline{\delta}$, such as $\overline{\delta}=1$. Then they can select an empirically plausible value of $M$, the largest allowed value of the omitted variable bias. Then remove any elements from ¯^$\widehat{\overline{\mathcal{B}}}_{I}(\overline{\delta},R_{\text{long}}^{2})$ that lay outside the range $[\widehat{\beta}_{\text{med}}-M,\widehat{\beta}_{\text{med}}+M]$, because these are deemed a priori implausible. The set of all remaining elements is then the estimated set of long regression coefficient values consistent with $|\delta|\leq\overline{\delta}$, the choice of $R_{\text{long}}^{2}$, and the restriction that the magnitude of OVB is no larger than $M$. This set can be interpreted as a set of bias corrected coefficients. For example, consider the specification in the second row of Table 6, with $M=2|\widehat{\beta}_{\text{med}}|=2\times 0.17=0.34$ (column (6)). In this case, ¯^$\widehat{\overline{\mathcal{B}}}_{I}(1,1.3\widehat{R}_{\text{med}}^{2})\cap[-0.17,0.51]=\big{(}(-\infty,-0.84]\cup[0.17,0.19]\big{)}\cap[-0.17,0.51]=[0.17,0.19]$. Thus researchers can state:
    “Assuming the omitted variable bias is no larger than $0.34$, and that selection on unobservables is no larger than selection on observables, the bias corrected set of coefficient estimates is $[0.17,0.19]$.”
  2. 2.
    Second, in lieu of computing explain away breakdown points, researchers can estimate the sign change breakdown point for an empirically plausible value of $M$. For example, consider again the specification in the second row of table 6, with $M=2|\widehat{\beta}_{\text{med}}|=0.34$ (column (6)). Here the estimated sign change breakdown point is 2.83. Thus researchers can state:
    “Assuming the omitted variable bias is no larger than 0.34, we estimate that selection on unobservables would need to be at least 2.83 times as large as selection on observables to change the sign of our baseline result.”
Whenever summary statements like these are made, we recommend that researchers discuss their choice of $M$, report their choice of $R_{\text{long}}^{2}$ (and consider exploring additional choices, see Section 5.3), and also still present identified set plots like in Figure 2, if only in an appendix.
@@ 7.1 Data
The second meta-analysis is based on papers used in the meta-analysis of Oster (2019, sections 2 and 5). This dataset has 141 regressions from 55 different papers that were published in either a top 5 economics journal or the AEJ: Applied Economics between 2008 and 2013, and which reported results from coefficient stability analyses. See Oster (2019) for more details on the construction of this dataset. Since this dataset already only contains only a handful of regressions per paper, we did not further restrict the sample. This second dataset is useful both because it provides a larger sample of regressions and also because the inclusion criteria are qualitatively different, since none of these papers implemented Oster’s methods because they preceded her paper.
@@ 7.2 Results
First consider the sample in Panel A. As shown in Theorem 2, the sign change breakdown point without any magnitude restrictions is never larger than 1. This result can be seen in the last line of all panels. For the rule of thumb choice of $R_{\text{long}}^{2}$, most of these breakdown points exactly equal 1. For $R_{\text{long}}^{2}=1$, more than half of the regressions are no longer robust when using $1$ as the robustness threshold point. In practice, researchers almost always go beyond binary robust / not-robust conclusions and also report the magnitude of the explain away breakdown point, interpreting larger magnitudes as evidence of greater robustness. The second row of each sub-panel shows that even with a reasonably strong restriction on the magnitude of OVB, the sign change breakdown points are substantially smaller than the explain away breakdown points. For example, the median is cut in half. Near the top of the distribution, the 90th percentile shrinks 87% from 13.22 down to 1.66 for the rule of thumb $R_{\text{long}}^{2}$, and similarly for $R_{\text{long}}^{2}=1$. The results for Panel B are qualitatively similar: The sign change breakdown points are substantially smaller in magnitude than the explain away breakdown points. Overall, these results show that in most empirical applications, common conclusions like
“selection on unobservables would have to be at least 4.1 times stronger than selection on observables to explain away the relationship” (Squicciarini 2020, AER, page 20) [Note: This computation uses $R_{\text{long}}^{2}=1$]
@@
$R^{2}_{\text{long}}$ 10% 25% 50% 75% 90%
@@
$\min\{1.3\widehat{R}_{\text{med}}^{2},1\}$ 1.0 1.0 1.0 1.3 67.2
1 1.0 1.0 1.0 13.8 21.3
@@
  • Note: Both samples only include results considered robust according to standard practice (the explain away breakdown point is weakly larger than 1, where we use the value of $R_{\text{long}}^{2}$ specified in the corresponding panel). Panel A: First row: $N=29$ regressions from 11 papers. Second row: $N=11$ regressions from 6 papers. Panel B: First row: $N=111$ regressions from 47 papers. Second row: $N=58$ regressions from 31 papers.
@@
To make magnitudes of OVB comparable across different regressions, we work with magnitudes relative to the baseline estimates $\widehat{\beta}_{\text{med}}$. Specifically, for the $i$th regression, let $\widehat{p}_{i}^{\text{bp}}$ be the largest magnitude of OVB (defined as a proportion of the baseline estimate) such that the estimated sign change breakdown point is at least as large as the estimated explain away breakdown point in this regression. [^4] Formally, we define $\widehat{p}_{i}^{\text{bp}}\coloneqq\sup\{p\geq 1:|\widehat{\delta}_{i}^{\text{bp},\text{sign}}(R_{\text{long}}^{2},p\cdot|\widehat{\beta}_{\text{med},i}|)|\geq|\widehat{\delta}^{\text{bp},\text{explain away}}_{i}(R_{\text{long}}^{2})|\}$. The two panels in Table 7.2 give the distribution of $\widehat{p}_{i}^{\text{bp}}$ across regressions $i$ in our meta-analysis samples for different choices of $R_{\text{long}}^{2}$.
@@
Here we see that, across both samples and both choices of $R_{\text{long}}^{2}$, more than 50% of regressions require that the sign of the parameter of interest $\beta_{\text{long}}$ must be known a priori if we want to interpret the explain away breakdown point as a sign change breakdown point. That follows since—recall from Section 3.2—the choice $M=\beta_{\text{med}}$ (equivalently, $p=1$) implies that the magnitude restriction A5 is equivalent to $\beta_{\text{long}}\in[0,2\beta_{\text{med}}]$ (when the baseline estimand is positive; the negative case is analogous). If the sign of $\beta_{\text{long}}$ was known a priori, however, then there would be no need to perform a sensitivity analysis to assess whether omitted variables could flip the sign of one’s baseline results; they cannot by assumption.
@@
For regressions above the 50th percentile, the implicit magnitude restriction is typically larger than $p=1$, but is often still quite restrictive. For example, the 75th percentile in Panel A is 1.3 under the rule of thumb $R_{\text{long}}^{2}$ choice. To interpret this, suppose $\widehat{\beta}_{\text{med}}=2$. Then $p=1.3$ corresponds to the assumption that $\beta_{\text{long}}\in[-0.6,4.6]$. Additional related results for these meta-analyses are in Appendix LABEL:sec:extra_meta-analyses.